OpenStudy (anonymous):

Given the system of equations presented here: 2x + 4y = 14 4x + y = 20 Which of the following actions creates an equivalent system such that, when combined with the other equation, one of the variables is eliminated?

pooja195 (pooja195):

Following....

OpenStudy (anonymous):

huh

pooja195 (pooja195):

"Which of the following actions creates an equivalent system such that, when combined with the other equation, one of the variables is eliminated? " Follwoing...

OpenStudy (anonymous):

Multiply the second equation by −4 to get −16x − 4y = −80 Multiply the second equation by −1 to get −4x − y = −20 Multiply the first equation by 2 to get 4x + 8y = 28 Multiply the first equation by −1 to get −2x − 4y = −14

pooja195 (pooja195):

Well opposites cancel So a negative # and a positive # cancel out What do you think that answer is?

OpenStudy (anonymous):

A?

pooja195 (pooja195):

Thats correct :)